Red Shoe Co. has concluded that additional equity financing will be needed to expand operations and that the needed funds will be best obtained through a rights offering. It has correctly determined that as a result of the rights offering, the share price will fall from $53 to $50.00 ($53 is the rights-on price: $50.00 is the ex-rights price, also known as the when-issued price). The company is seeking $13 million in additional funds with a per-share subscription price equal to $40. 


How many shares are there currently, before the offering? (Assume that the increment to the market value of the equity equals the gross proceeds from the offering.) (Do not round intermediate calculations and round your answer to nearest whole number, e.g., 32.) 


Number of old shares = _______ 
Answer #1

The number of old shares is computed as shown below:

Number of shares a shareholder will have ex rights is computed as follows:

Ex right price = [ N ( $ 53 ) + $ 40 ] / ( N + 1 )

$ 50 = [ N ( $ 53 ) + $ 40 ] / ( N + 1 )

N = 3.333333333

Number of new shares is computed as follows:

= Additional funding / Per share subscription price

= $ 13,000,000 / $ 40

= 325,000

So the number of old shares will be:

= 3.333333333 x 325,000

= 1,083,333 Approximately

Feel free to ask in case of any query relating to this question
Answer #2

current market price(p0)= $53

Ex rights price(pe)= $ 50

subscription price(ps)= $40

New no of shares = funds to be raised/ ps = 13000000/40 = 325000 stocks

Also, pe= p0-value of a right(vr) 

So, vr= p0 - pe= 53-50 = $3

Again, Vr= (p0-ps)/(*+1), where *= value of a right

so *= 3.33 rights

finally, *= old shares/ new shares

therefore old shares= 325000* 3.33= 1083333.33